# MedIAnomaly: A comparative study of anomaly detection in medical images

### v2

@@ 1 Introduction
Although some surveys and benchmarks have been made for medical AD [35, 7, 57, 5, 23], they do not provide a comprehensive and fair comparison of methods. Fernando et al. [35] provided a systematic review of machine learning–based medical AD techniques with their applications and limitations, while no experiments were conducted. Baur et al. [7] compared a series of medical reconstruction-based AD methods on unified datasets. However, the comparison was limited to reconstruction methods only and the experiments were performed on a single image modality (brain MRI). These could lead to incomplete conclusions. Also, their utilization of private datasets impedes reproducibility. Recently, Bao et al. [5] and Lagogiannis et al. [57] evaluated various AD methods on reorganized public medical datasets. However, most of the methods included by Bao et al. [5] were originally proposed for industrial scenes instead of the medical domain. While Lagogiannis et al. [57] evaluated various SOTA medical AD methods, as well as several best-performing reconstruction-based methods from Baur et al. [7], they did not delve into detailed analysis of those typical methods. Furthermore, they did not study the effect of network architectures. Cai et al. [23] evaluated various medical AD methods on reorganized datasets, but the effect of architectures was still unexplored and the methods involved were not comprehensive.
To address the aforementioned issues, we aim to present a comprehensive and fair comparison for typical AD methods on unified medical benchmarks. The main contribution of this work is summarized as follows:
@@
    We conduct a comprehensive and unified comparison of thirty representative methods using the collected datasets. Furthermore, the effects of key components employed in these methods are analyzed.
@@ 2.2.3 Other SSL-related approaches
It is noteworthy that many AD methods in other taxonomies also incorporate SSL designs to improve their performance. However, we do not categorize these methods as SSL-based AD methods due to their distinct working principles, which differ from our definition of SSL-based AD methods but align with other taxonomies. For example, Zimmerer et al. [127] introduced an inpainting task to AE, and Kascenas et al. [51, 52] introduced a denoising task to UNet. Although these methods leverage self-supervised tasks during training, their objective is to facilitate pseudo-healthy reconstruction, which is the goal of reconstruction-based AD methods. Therefore, these methods are considered reconstruction-based methods rather than SSL-based methods.
@@ 3.1.4 Summary of the datasets
Table 2: Summary of the datasets.
Dataset Modality Task Repartition
$\mathcal{D}_{train}$ $\mathcal{D}_{test}$ (Normal+Abnormal)
RSNA2 Chest X-ray AnoCls 3851 1000+1000
VinDr-CXR3 [73] Chest X-ray AnoCls 4000 1000+1000
Brain Tumor4 Brain MRI AnoCls 1000 600+600
LAG [60] Retinal fundus image AnoCls 1500 811+811
ISIC2018 [33] Dermatoscopic image AnoCls 6705 909+603
Camelyon16 [11] Histopathology image AnoCls 5088 1120+1113
BraTS2021 [3] Brain MRI AnoCls & AnoSeg 4211 828+1948
@@
It is noteworthy that the datasets display distinct abnormal patterns. The chest X-ray, brain MRI, and retinal fundus datasets (i.e., RSNA, VinDr-CXR, Brain Tumor, BraTS2021 and LAG) showcase local anomalies. This implies that abnormal images contain the healthy version, with the addition of lesions in some areas. In contrast, Camelyon16 and ISIC2018 present global semantic anomalies, wherein abnormal images belong to completely different categories compared to normal images.
@@ 4 Experiments and analysis
AnoDDPM [113] [Reference to the original paper]
AutoDDPM [12] [Reference to the original paper]
@@
To provide a comprehensive evaluation and analysis for medical AD methods, we assess a total of thirty AD methods on five different image modalities. This extensive coverage guarantees the generalizability of our results and findings. Table 5 provides a summary of the methods in our experiments.
@@ 4.1 Comparison of the state-of-the-arts
For pixel-level AnoSeg, reconstruction methods significantly outperform SSL methods . As shown in Table 7, the vanilla AE outperforms all SSL one-stage segmentation methods in terms of the segmentation metrics on BraTS2021. This can be attributed to the fact that SSL methods trained through synthetic data segmentation rely on the similarity between synthetic abnormal patterns and real anomalies. They struggle to recognize real abnormal regions when these regions exhibit significant differences in appearance compared to synthetic anomalies. Consequently, their performance deteriorates on such datasets. This point will be further discussed in Section 4.3.1.
@@
Fig. 5: Visualization of the compared image-reconstruction methods on typical abnormal images. The first row displays input abnormal images along with the abnormal regions (if available). The subsequent rows depict the reconstructed images and anomaly maps generated by each method.
@@ 4.2.2 Effects of latent space restrictions in AE
In addition to the simplest latent dimension reduction, there exist other popular designs that impose latent space restrictions in various ways, such as VAE [54], Constrained AE [29], and MemAE [36]. However, we observe from Table 6 that none of the existing approaches for restricting the latent space demonstrate superior performance compared to latent dimension reduction. Potential reasons are that VAE samples the latent code from the learnt distribution instead of using a deterministic vector, making the reconstruction indeterministic, while Constrained AE does not explicitly restrict the latent space.
The unexpected failure of MemAE [36], despite its success on various natural images, raises questions. MemAE introduced a memory bank to store prototypes of normal latent features during training and utilized a combination of these prototypes to replace the latent feature of the input. It hypothesized a distinct difference between normal and abnormal features, aiming to render abnormal inputs unreconstructible using the combination of normal prototypes. However, this hypothesis is not necessarily valid due to the inherent nature of AE [13, 21]. Consequently, we attribute the failure of MemAE to the inseparability of latent features of normal and abnormal medical images. To substantiate our argument, we visualize the latent features of normal and abnormal medical images obtained from AE and present the performance of OC-SVM, which is built on these features, in Fig. 7. The results clearly demonstrate that these features are mixed together and inseparable by OC-SVM. This finding contradicts the underlying hypothesis of MemAE and provides support for our argument, ultimately explaining the failure of MemAE in medical scenarios.
@@ 4.2.6 Reconstruction quality and AD performance
Previously, Baur et al. [7] demonstrated that there is a relatively weak correlation between reconstruction quality and AnoSeg performance specifically on brain MRIs. They observed that in certain cases, as the reconstruction error increases (i.e., the reconstruction quality decreases), the AnoSeg performance actually improves. Building upon their findings and conducting experiments on a wider range of datasets and methods, we extend their conclusion to encompass additional data modalities and provide further insights to certain datasets.
@@
Expanding on the conclusion drawn by Baur et al. [7], we showcase that the relationship between reconstruction quality and AD performance is not only applicable to brain MRIs but also extends to other modalities such as chest X-rays, retinal fundus images, and dermatoscopic images. A prominent example supporting this conclusion is the comparison between AnoDDPMGaussian and AE-PL. As depicted in Fig. 5, AnoDDPMGaussian achieves the highest reconstruction quality on all the datasets, effectively preserving both normal and abnormal regions. However, as highlighted in Tables 6 and 7, AnoDDPMGaussian exhibits the poorest overall performance. This can be attributed to the ineffectiveness of Gaussian noise in corrupting low-frequency abnormal regions, leading to their well-reconstruction alongside normal regions. Conversely, despite generating visually “faulty” reconstructions, AE-PL demonstrates competitive performance across most datasets. The potential reason is that the perceptual loss aids in detecting abnormal regions in the feature space. Therefore, the emphasis is placed on capturing high-level semantic information rather than intensity reconstruction.
@@
While our results reinforce the conclusion made by Baur et al. [7] that high reconstruction quality is usually unnecessary, we also uncover a complementary finding regarding the detection of specific abnormal patterns, such as tumor cells in histopathology images (Camelyon16), where high reconstruction quality can yield benefits. Combining Table 6 and Fig. 5, we observe that, excluding methods that leverage ImageNet weights, the top three performing methods on Camelyon16 are AutoDDPM, AnoDDPMSimplex, and DAE. Notably, these methods also generate high-quality reconstructions. We attribute this observation to the distinctive abnormal patterns present in histopathology images, where tumor cells exhibit unexpected morphological variations compared to normal cells. To effectively recognize these anomalies related to cellular morphology in the image space, high reconstruction quality proves to be beneficial.
@@ 5.2 Distance function for measuring reconstruction error
The concept of learning distance functions (similarity metrics) through deep learning was originally proposed for multimodal medical image registration. Traditional metrics often struggle to measure the similarity between different imaging modalities due to the high variability of tissue appearance. Simonovsky et al. [101] addressed this challenge by training a network to measure similarity using constructed image pairs. Similarly, in anomaly detection, certain reconstruction errors are unrelated to anomalies and can be attributed to limited model capacity. These errors should ideally be disregarded when calculating the distance between the input and the reconstruction, with emphasis placed solely on highlighting abnormal regions. Drawing inspiration from deep similarity metrics in registration, we believe it is possible to train a similarity network to automatically measure anomaly-related reconstruction errors. This network can be trained on synthesized image pairs generated from the normal training set, taking into account the data characteristics. Pursuing this direction may lead to the development of a robust distance function capable of effectively measuring reconstruction errors across various datasets with different image modalities.
@@ 5.3 Utilization of ImageNet pre-trained weights
While ImageNet weights have been effectively applied in various contexts, it is important to highlight that their application in feature reference-based methods, such as knowledge distillation [120] and feature modeling [58] approaches, which are particularly popular in the industrial domain, is not suitable for medical AD. As shown in Table 13, the performance of these methods in medical images is inferior to that of other strategies utilizing ImageNet weights, as indicated in Table 6. This observation prompts us to reconsider how we can enhance these commonly used strategies in industrial AD and adapt them effectively to the medical domain.

### v3

@@ 1 Introduction
Although some surveys and benchmarks have been made for medical AD [35, 7, 57, 5, 23], they do not provide a comprehensive and fair comparison of methods. Fernando et al. [35] provided a systematic review of machine learning–based medical AD techniques with their applications and limitations, while no experiments were conducted. Baur et al. [7] compared a series of medical reconstruction-based AD methods on unified datasets. However, the comparison was limited to reconstruction methods only and the experiments were performed on a single image modality (brain MRI). These could lead to incomplete conclusions. Also, their utilization of private datasets impedes reproducibility. Cai et al. [23] assessed various medical AD methods on reorganized datasets, but did not dive into a detailed analysis of these methods. Bao et al. [5] evaluated a range of AD methods on reorganized public medical datasets. However, most of these methods were originally designed for industrial applications rather than the medical domain.
It is important to note that Lagogiannis et al. [57] recently offered a detailed analysis of diverse SOTA AD methods across multiple datasets. They extensively investigated the influence of anomaly size and intensity, the effects of limited training data, and the performance of self-supervised pre-training on weight initialization and backbone pre-training. However, their analysis did not thoroughly encompass some non-latest yet representative methods, such as various variants of AE-based methods and certain self-supervised methods utilizing specialized synthetic anomalies. Additionally, they did not employ unified network architectures for analyzing methods within the same paradigm. As a result, the inherent properties of the components and network architectures in typical methods remain unexplored. Unlike Lagogiannis et al. [57], we aim to further investigate the effects and inherent properties of key components in typical methods by conducting comparisons across comprehensive datasets and employing as fair as possible network configurations. The main contribution of this work is summarized as follows:
@@
    We conduct a comprehensive comparison of thirty representative methods using the collected datasets. Furthermore, the effects of key components employed in these methods are analyzed.
@@ 2.2.3 Other SSL-related approaches
It is noteworthy that many AD methods in other taxonomies also incorporate SSL designs to improve their performance. However, we do not categorize these methods as SSL-based AD methods due to their distinct working principles, which differ from our definition of SSL-based AD methods but align with other taxonomies. For example, Zimmerer et al. [129] introduced an inpainting task to AE, and Kascenas et al. [51, 52] introduced a denoising task to UNet. Although these methods leverage self-supervised tasks during training, their objective is to facilitate pseudo-healthy reconstruction, which is the goal of reconstruction-based AD methods. Therefore, we categorize these methods as reconstruction-based rather than SSL-based methods [^2] Note that DAE [52] is also considered a SSL-based AD methods by some studies [57], as it trains a UNet with skip connections using a self-supervised denoising task. The categorization of DAE may still be controversial..
@@ 3.1.1 Datasets not suitable for anomaly detection
Table 2: Summary of the datasets.
Dataset Modality Task Repartition
$\mathcal{D}_{train}$ $\mathcal{D}_{test}$ (Normal+Abnormal)
RSNA3 Chest X-ray AnoCls 3851 1000+1000
VinDr-CXR4 [75] Chest X-ray AnoCls 4000 1000+1000
Brain Tumor5 Brain MRI AnoCls 1000 600+600
LAG [60] Retinal fundus image AnoCls 1500 811+811
ISIC2018 [33] Dermatoscopic image AnoCls 6705 909+603
Camelyon16 [11] Histopathology image AnoCls 5088 1120+1113
BraTS2021 [3] Brain MRI AnoCls & AnoSeg 4211 828+1948
@@ 3.1.3 Pixel-level anomaly segmentation datasets
(a) Train Normal
(b) Test Normal
(c) Test Abnormal
Fig. 4: Statistics of slice indices in our processed BraTS2021 dataset. A central crop was performed before slice extraction, resulting in index 0 corresponding to index 50 of the original volume and index 69 corresponding to index 119 of the original volume.
@@
Due to the intrinsic characteristics of brain MRIs, central slices typically exhibit more relevant content and anomalies, as illustrated in Fig. 4. This imbalance could introduce biases to slice-level anomaly classification. Consequently, this dataset is primarily intended for evaluating pixel-level anomaly segmentation, while we also provide image-level metrics for reference.
@@ 3.1.4 Summary of the datasets
The datasets display distinct abnormal patterns. The chest X-ray, brain MRI, and retinal fundus datasets (i.e., RSNA, VinDr-CXR, Brain Tumor, BraTS2021 and LAG) showcase local anomalies. This implies that abnormal images contain the healthy version, with the addition of lesions in some areas. In contrast, Camelyon16 and ISIC2018 present global semantic anomalies, wherein abnormal images belong to completely different categories compared to normal images.
@@ 4 Experiments and analysis
Intensity $\mathbf{x}_{i}$ or $-\mathbf{x}_{i}$ (subject to the dataset) $-\mathbf{x}_{i}$ for the Brain Tumor dataset $\mathbf{x}_{i}$ for others
@@
AnoDDPM [115] UNet [DDPM] $\mathbb{E}_{t\sim[1,T],x\sim q(x)}\left[\left\|\epsilon-\epsilon_{\theta}\left(x_{t},t\right)\right\|^{2}\right]$ $(\mathbf{x}_{i}-\hat{\mathbf{x}}_{i})^{2}$ $\epsilon$: Simplex noise $\epsilon_{\theta}$: UNet t: Time step Input size: $256\times 256$
AutoDDPM [12] UNet [DDPM] $\mathbb{E}_{t\sim[1,T],x\sim q(x),\epsilon\sim\mathcal{N}(0,1)}\left[\left\|\epsilon-\epsilon_{\theta}\left(x_{t},t\right)\right\|^{2}\right]$ $|\hat{\mathbf{x}}-\mathbf{x}|*\hat{m}$ Customized inference step: Mask, Stitch, and Re-Sample. $\hat{m}$: The initial mask
@@
To provide a comprehensive evaluation and analysis for medical AD methods, we assess a total of thirty AD methods on five different image modalities. This extensive coverage guarantees the generalizability of our results and findings. Table 5 provides a summary of the methods in our experiments. In addition, we evaluate the performance of directly using pixel intensity as the anomaly score [71], denoted as ‘Intensity’, which reflects the distinguishability between normal and abnormal pixel intensity within datasets.
@@ 4.1 Comparison of the state-of-the-arts
For pixel-level AnoSeg, reconstruction methods significantly outperform SSL methods. As shown in Table 7, the vanilla AE beats all SSL one-stage segmentation methods in terms of the segmentation metrics on BraTS2021. This can be attributed to the fact that SSL methods trained through synthetic data segmentation rely on the similarity between synthetic abnormal patterns and real anomalies. They struggle to recognize real abnormal regions when these regions exhibit significant differences in appearance compared to synthetic anomalies. Consequently, their performance deteriorates on such datasets. This point will be further discussed in Section 4.3.1.
@@
[…]
@@ 4.2.2 Effects of latent space restrictions in AE
In addition to the simplest latent dimension reduction, there exist other popular designs that impose latent space restrictions in various ways, such as VAE [54], Constrained AE [29], and MemAE [36]. However, we observe from Table 6 that none of the existing approaches for restricting the latent space demonstrate superior performance compared to latent dimension reduction. Potential reasons are that VAE samples the latent code from the learnt distribution instead of using a deterministic vector, making the reconstruction indeterministic, while Constrained AE does not explicitly restrict the latent space. The unexpected failure of MemAE [36], despite its success on various natural images, raises questions. MemAE introduced a memory bank to store prototypes of normal latent features during training and utilized a combination of these prototypes to replace the latent feature of the input. It hypothesized a distinct difference between normal and abnormal features, aiming to render abnormal inputs unreconstructible using the combination of normal prototypes. However, this hypothesis is not necessarily valid due to the inherent nature of AE [13, 21]. Consequently, we attribute the failure of MemAE to the inseparability of latent features of normal and abnormal medical images. To substantiate our argument, we visualize the latent features of normal and abnormal medical images obtained from AE and present the performance of OC-SVM, which is built on these features, in Fig. 8. The results clearly demonstrate that these features are mixed together and inseparable by OC-SVM. This finding contradicts the underlying hypothesis of MemAE and provides support for our argument, ultimately explaining the failure of MemAE in medical scenarios.
@@ 4.2.6 Reconstruction quality and AD performance
Correlation. Previously, Baur et al. [7] demonstrated a relatively weak correlation between reconstruction quality and AnoSeg performance on brain MRIs. Instead, it is the distinguishable residual histograms of normal and anomalous intensities that matter most. They observed that in certain cases, as the reconstruction error increases (i.e., the reconstruction quality decreases), the AnoSeg performance actually improves. Building upon their findings and conducting experiments on a wider range of datasets and methods, we extend their conclusion to encompass additional data modalities and provide further insights to certain datasets.
@@
Expanding on the conclusion drawn by Baur et al. [7], we show that the relatively weak correlation between reconstruction quality and AD performance is also applicable to other modalities such as chest X-rays, retinal fundus images, and dermatoscopic images. A prominent example supporting this conclusion is the comparison between AnoDDPMGaussian and AE-PL. As depicted in Fig. 6, AnoDDPMGaussian achieves the highest reconstruction quality on all datasets, effectively preserving both normal and abnormal regions. However, as highlighted in Tables 6 and 7, AnoDDPMGaussian exhibits the poorest overall performance. This can be attributed to the ineffectiveness of Gaussian noise in corrupting low-frequency abnormal regions, leading to their well-reconstruction alongside normal regions. Conversely, despite generating visually “faulty” reconstructions, AE-PL demonstrates competitive performance across most datasets. The reason is that perceptual loss performs anomaly detection in the feature space, prioritizing high-level semantic information over intensity reconstruction.
It is noteworthy that some recent works [12, 62] achieve both high reconstruction quality and competitive AD performance. These studies concentrate on enhancing the reconstruction quality of normal areas, rather than the entire image, and employ strategies such as masking to degrade the reconstruction quality of abnormal regions. Therefore, they still support the conclusion that distinguishable residual histograms of normal and anomalous intensities are more crucial than the overall reconstruction quality.
@@
Additionally, we reveal a complementary insight concerning the detection of specific abnormal patterns, such as tumor cells in histopathology images (Camelyon16), where high reconstruction quality can yield benefits. Combining Table 6 and Fig. 6, we observe that, excluding methods equipped with ImageNet weights, the top three performing methods on Camelyon16 – AutoDDPM, AnoDDPMSimplex, and DAE – consistently produce high-fidelity reconstructions. We attribute this observation to the distinctive abnormal patterns present in histopathology images, where tumor cells manifest morphological deviations compared to normal cells. High reconstruction quality is beneficial for recognizing these anomalies related to cellular morphology.
Fig. 9: Distribution of pixel-level anomaly scores calculated in image-space and feature-space. Scores are normalized through division by the maximum value. Note the substantial difference in the vertical scale of (a) and (b).
Characteristics of relevant methods. To investigate the disparity between reconstruction quality and performance metrics, we began from analyzing the binary prediction maps depicted in Fig. 6 (the third column of BraTS2021 dataset). The binary prediction maps illustrate that methods such as DAE and autoDDPM, known for their high-quality reconstruction, excel in segmenting small tumors. However, their prediction maps contain scattered noisy points on normal regions. In contrast, methods such as AE-PL, which generate coarse score maps from feature-space, tend to produce larger segmentation regions than actual abnormal areas but exhibit improved robustness in normal regions without introducing noisy points. These characteristics could be relevant to the calculation space (i.e., image-space and feature-space) of their anomaly scores.
For further examination, we visualized the distribution of pixel-level anomaly scores for typical methods in Fig. 9. Initially, Fig. 9 shows that image-space anomaly scores are concentrated around zero, reflecting high reconstruction quality of corresponding methods. Improvements in these methods should prioritize amplifying errors in abnormal regions to reduce indistinguishable anomalies. For instance, autoDDPM exhibits notably lower frequencies of abnormal histogram overlapping with normal ones near zero compared to AnoDDPM (Gaussian), consistent with its superior performance metrics. In contrast to the lower values of image-space anomaly scores, feature-space scores from AE-PL and FAE exhibit values distant from zero on both normal and abnormal pixels, with distinguishable scores between the two, explaining their coarse score maps yet high performance metrics.
Analyzing these characteristics is crucial for clinical practice. However, existing quantitative metrics may not fully capture their distinctions to reflect clinical utility. For instance, widely used metrics like Dice score favor overconfident predictions, thus benefiting methods inclined towards over-segmentation like AE-PL and FAE. Furthermore, while feature-space anomaly scores from FAE and AE-PL distinguish between normal and abnormal pixels and achieve satisfactory performance metrics, they pose a challenge in clinical scenarios where discerning intensity differences in non-zero positive scores without proper calibration can be difficult for human experts. Future efforts should develop metrics and calibration strategies to effectively evaluate and enhance the clinical utility of anomaly segmentation.
@@ 5 Challenges and future directions
Fig. 11: Intensity distribution of normal and abnormal pixels in BraTS2021 dataset.
5.1 Anomalies with pronounced intensity differences
[…]
The current analysis of pixel-level AnoSeg remains limited due to the scarcity of annotations and the narrow range of anomaly types, primarily focusing on prominent tumors in brain MRIs. Future work should expand the evaluation to include diverse and subtle anomalies, which could better reveal the potential of deep learning methods and provide deeper insights into the capabilities and limitations of AD methods in real-world clinical scenarios.
@@ 5.3 Distance function for measuring reconstruction error
The concept of learning distance functions (similarity metrics) through deep learning was originally proposed for multimodal medical image registration. Traditional metrics often struggle to measure the similarity between different imaging modalities due to the high variability of tissue appearance. Simonovsky et al. [103] addressed this challenge by training a network to measure similarity using constructed image pairs. Analogously, in anomaly detection, certain reconstruction discrepancies may not necessarily correspond to actual anomalies but rather reflect inherent limitations in model capacity. Ideally, such discrepancies should be systematically excluded when computing the distance metric between input and reconstructed images, with emphasis placed solely on highlighting abnormal regions. Drawing inspiration from deep similarity metrics in registration, we believe it is possible to train a similarity network to automatically measure anomaly-related reconstruction errors. This network can be trained on synthesized image pairs generated from the normal training set, taking into account the data characteristics. Pursuing this direction may lead to the development of a robust distance function capable of effectively measuring reconstruction errors across various datasets with different image modalities.
@@ 5.4 Utilization of ImageNet pre-trained weights
While ImageNet weights have been effectively applied in various contexts, it is important to highlight that their application in feature reference-based methods, such as knowledge distillation [122] and feature modeling [58] approaches, which are particularly popular in the industrial domain, is not suitable for medical AD. The performance of these methods in medical images, as shown in Table 13, is inferior to that of other strategies utilizing ImageNet weights, as indicated in Table 6. This observation prompts us to reconsider how we can enhance these commonly used strategies in industrial AD and adapt them effectively to the medical domain.